# eMoE-Tracker: Environmental MoE-based Transformer for Robust Event-guided Object Tracking

### v2

@@ Abstract
The unique complementarity of frame-based and event cameras for high frame rate object tracking has recently inspired some research attempts to develop multi-modal fusion approaches. However, these methods directly fuse both modalities and thus ignore the environmental attributes, e.g., motion blur, illumination variance, occlusion, scale variation, etc. Meanwhile, no interaction between search and template features makes distinguishing target objects and backgrounds difficult. As a result, performance degradation is induced especially in challenging conditions. This paper proposes a novel and effective Transformer-based event-guided tracking framework, called eMoE-Tracker, which achieves new SOTA performance under various conditions. Our key idea is to disentangle the environment into several learnable attributes to dynamically learn the attribute-specific features and strengthen the target information by improving the interaction between the target template and search regions. To achieve the goal, we first propose an environmental Mix-of-Experts (eMoE) module that is built upon the environmental Attributes Disentanglement to learn attribute-specific features and environmental Attributes Gating to assemble the attribute-specific features by the learnable attribute scores dynamically. The eMoE module is a subtle router that fine-tunes the transformer backbone more efficiently. We then introduce a contrastive relation modeling (CRM) module to emphasize target information by leveraging a contrastive learning strategy between the target template and search regions. Extensive experiments on diverse event-based benchmark datasets showcase the superior performance of our eMoE-Tracker compared to the prior arts. Project page: https://vlislab22.github.io/eMoE-Tracker/
@@
Due to the limited space in the main paper, we provide additional material for the proposed method and experimental results. Sec. VI introduce the datasets. Then, Sec. VII illustrates more details about the implementation and experiments. Afterward, we report more visual results and performance evaluation under different attributes in Sec. VIII. In the end, Sec. IX makes summary for the complementary material.
@@ I INTRODUCTION
Figure 1: An illustration of the core idea of the environmental MoE (eMoE) module. This module acts as a subtle router to fine-tune the frozen backbone encoder. The number of experts is determined by the attributes we decouple for the environmental conditions, and each expert is responsible for learning the attribute-specific features. All the learned features are assembled and added with the ones from the backbone encoder at the corresponding layer for robust tracking representation.
@@
This has inspired research endeavors in developing event-guided, i.e., RGB-event (RGB-E) multi-modal tracking approaches [4, 5, 6, 7, 8, 9, 10]. These works can be divided into two categories based on the network structure: two-stream, i.e., siamese network [5, 6, 7] and one-stream trackers [4, 9, 10]. The former takes two identical branches to process RGB and event modality separately. To better leverage the complementarity and increase the interaction between them, complex fusion modules are designed, thus leading to model complexity. The latter is usually based on the vision transformer (ViT) structure [11], where RGB and event tokens are concatenated and fed into the ViT backbone for feature encoding. Although they are free from the complex network structure, they fail to consider the impact of environmental attributes on tracking performance in challenging conditions. Meanwhile, no interaction between search and template features makes distinguishing target objects and backgrounds difficult. Consequently, performance degradation is induced especially in challenging conditions. Intuitively, we raise a novel research question: how to design a one-stream framework that can distinguish the environmental attributes while enabling feature interaction for robust tracking under diverse visual conditions?
@@
In this paper, we propose a novel one-stream framework with an environmental Mixture-of-Experts structure (eMoE) along with a contrastive relation modeling (CRM) module to achieve robust tracking in challenging conditions, as shown in Fig. 1. The key insight is to disentangle the environmental attributes through learnable layers to dynamically learn the attribute-specific features for better interaction between the target objects and background. Specifically, to disentangle the environmental attributes, the eMoE module ((Sec. III-B)) is proposed to achieve two goals: (i) the environmental attributes disentanglement and (ii) the environmental attributes assembling. For the former, the eMoE module disentangles four attributes –illumination variance, motion blur, scale variance, and occlusion – to learn the attribute-specific features. This has been experimentally shown sufficient to reflect environmental effects on tracking given the advantages of event cameras (See Tab. IV). For the latter, each attribute-specific feature is assembled to build a more discriminative representation for tracking w.r.t.the attribute scores under the corresponding visual conditions, e.g., motion blur. For more efficient training, our eMoE module can be inserted into the arbitrary layers to fine-tune the ViT backbone encoder. For the proposed CRM (Sec. III-B3) module, it aims to better distinguish the target features and background ones by introducing contrastive learning strategy. This subtly improves the interaction between the target template and search region and enhances the target objects. By integrating the eMoE and CRM modules, the output features can be more discriminative and less noisy for more robust tracking performance under diverse visual conditions.
@@ III Method
Figure 2: Overview of our proposed framework. The input of the whole network is the patch embeddings of RGB frames and stacked event frames. The concatenated two modal patches are fed into the backbone model and eMoE, and eMoE is inserted to $L$-th ViT to generate feature tokens which are combined with the tokens from the ViT encoder at the corresponding layer. $E^{l}$ is the ViT encoder at layer $l$. The CRM module gets the enhanced tokens to further improve the discriminability of target object.
@@ III-A1 Problem Setting
Given an initial target bounding box $B_{0}$ in a video, the goal of an RGB-based tracker is to learn a tracking model $T_{RGB}:\{I_{RGB},B_{0}\}\rightarrow B$ to estimate the bounding box in all subsequent frames $I_{RGB}$. In RGB-E tracking, event streams are introduced and stacked as event frames, extending the input to $(I_{RGB},I_{E})$, where the subscript $E$ indicates events. For details of representations for event data, we refer readers to [3, 29]. Therefore, the RGB-E tracking model can be represented as $T_{RGB-E}:\{I_{RGB},I_{E},B_{0}\}\rightarrow B$. We choose the transformer encoder and decoder of  [4] as our backbone encoder and decoder for better efficiency. The archetypical structure of the backbone encoder and decoder can be represented as $\mathcal{F}_{E}\circ\mathcal{F}_{D}$, where $\mathcal{F}_{E}:\{I_{RGB},I_{E},B_{0}\}\rightarrow\mathcal{T}_{RGB-E}$ denotes the backbone encoder and $\mathcal{F}_{D}:\mathcal{T}_{RGB-E}\rightarrow B$ represents the decoder which produces the estimated bounding box results. The main body of $\mathcal{F}_{D}$ here is a vanilla vision transformer [11] containing 12 encoder layers. Each layer contains Multi-head Self-Attention (MSA), LayerNorm (LN), Feed-Forward Network (FFN) and residual connections. Before feeding inputs into the backbone network, RGB and event patches are projected into feature tokens adding with positional embeddings and then concatenated to RGB and event feature tokens $\mathcal{T}_{RGB-E}^{0}=[\mathcal{T}_{RGB}^{z},\mathcal{T}_{RGB}^{x},\mathcal{T}_{E}^{z},\mathcal{T}_{E}^{x}]$ as the inputs of the transformer encoder. Tokens through the $l$-th encoder layer $E^{l}$ can be represented as $\mathcal{T}_{RGB-E}^{l-1}$. The final layer encoder output is denoted as $\mathcal{T}_{RGB-E}^{L}$.
@@ III-B1 Overview
An overview of our eMoE-Tracker is shown in Fig. 2. The RGB and event inputs are first projected into a sequence of tokens and then fed into backbone encoders and eMoE. The eMoE module aims to achieve environmental attributes disentanglement and environmental attributes gating. The backbone encoder layers are frozen and the parameters are not updated. The eMoE module disentangles the environmental attributes to learn the attribute-specific features.
Outputs from the eMoE module can be dynamically added to the tokens from the corresponding layer of the ViT backbone. Overall, the process can be formulated as follows:
@@
The eMoE module aims to achieve: i) environmental attributes disentanglement and ii) environmental attributes gating. We now describe them.
i. Environmental Attributes Disentanglement. As shown in Fig. 3, to enable better learning of the attribute-specific features under various challenging conditions, we manually annotate the visible-event datasets with four attribute labels including motion blur, illumination variance, scale variance, and occlusion. Then, a mixture-of-experts network with four identical branches is designed to learn the attribute-specific features for each challenging scenario. It allows us to capture more discriminative features and suppress the noises brought by other environmental attributes. All four expert networks employ the CONV-MLP-CONV structure but with different parameters. Specifically, considering the $l$-th ViT layer, we assume that there are $K$ experts $\{f_{expert}^{l,i}(\mathcal{T}^{l}):\mathbb{R}^{N\times D}\rightarrow\mathbb{R}^{N\times D},i\in[1,K]\}$ to learn the attribute-specific features under the corresponding environmental condition, where $l$ denotes the layer of ViT backbone and $i$ represents the index of experts. Through each expert, we generate a series of attribute-specific features $\{\mathcal{H}_{i}^{l}\in\mathbb{R}^{N\times D},i\in[1,K]\}$ by expert $f_{expert}^{l,i}(\mathcal{T}^{l})$
ii. Environmental Attribute Assembling. After obtaining the attribute-specific features under different decoupled environmental conditions, we should consider the different contributions of these features with the supervision of the ground truth attribute labels $G=[G_{1},G_{2},...,G_{K}]$. To achieve the goal, the gating network employs the CONV-BatchNorm-ReLU-CONV-Sigmoid structure with two loops and is fed by all the RGB and event patch tokens to generate $K$ attribute scores for attribute-specific features, where $K$ denotes the number of experts. The learnable score indicates the main challenging types in the corresponding scenario, therefore the attribute-specific feature with a larger score should have a higher contribution to the assembling features to achieve the robust representation under various challenging conditions. Moreover, it can suppress the noise from other environmental attributes. Specifically, at $l$-th layer of the backbone, the attribute scores $W^{l,t}$ are generated from the gating network $\{f_{g}^{l,t}(\mathcal{T}^{l}):\mathbb{R}^{N\times D}\rightarrow\mathbb{R}^{K}\}$, where $t$ represent the index of experts. The assembling feature $\mathcal{F}_{assemble}^{l}$ at layer $l$ can be formally calculated by $\sum_{t=1}^{K}W^{l,t}\mathcal{H}_{t}^{l}.$
@@ III-B3 Contrastive Relation Modeling
Apart from increasing the distinguishable ability of representations in different environmental conditions, we want to enhance the target information by introducing more interaction between target template and search region. Based on the fact that the features from the target template primarily contains target information, while the search regions include both the target and background information, we propose a CRM module by leveraging the contrastive learning strategy. As shown in Fig. 4, we first fuse two-modal patch tokens into fused target template tokens and search region tokens to better build the relationship. After fusion, we create positive pairs between features from target template and target object contents of search regions, and negative pairs between features from target template and background contents of search regions. To better strengthen the target object information, contrastive learning mechanism is exploited to pull the positive pairs closer and push negative pairs away thereby allowing the target object and background more distinguishable unambiguous. The proposed CRM module effectively help generate more unambiguous representation and achieve high performance in various challenging conditions.
@@ III-C Optimization
The main body of our RGB-E tracker $F_{RGB-E}$ is initialized by the transformer-based tracking backbone. All the parameters $\theta$ we should update is only existing in eMoE and CRM. The optimization process can be formulated as
@@
where $L_{cls}$ is the focal loss [30] for classification, IoU loss [31] $L_{iou}$ and $L_{1}$ are exploited for bounding box regression, $\lambda_{iou}$ and $\lambda_{L_{1}}$ are regularization parameters. For more details,  [22] can be referred to. Additionally, we take the InfoNCE loss [32] as a contrastive learning loss for the CRM module. Given the fused target template tokens $\mathcal{T}_{fused}^{z}=[t_{fused}^{z,1},t_{fused}^{z,2},...,t_{fused}^{z,N_{z}}]$ from the backbone ViT encoder, we compute the similarity $S=[s^{1},s^{2},...,s^{N_{x}}]$ between $\mathcal{T}_{fused}^{z}$ and the fused search region tokens $\mathcal{T}_{fused}^{x}=[t_{fused}^{x,1},t_{fused}^{x,2},...,t_{fused}^{x,N_{x}}]$. where $\tau$ is the temperature parameter and $s^{i}=sim(\mathcal{T}_{fused}^{z},t_{fused}^{x,i})/\tau$. Based on that, the search region tokens which contain the information inside the ground-truth bounding box can be selected as positive and the similarity score is defined as $s_{p}$ and the left are negative pairs including $N_{neg}$, the similarity score is set as $[s_{n}^{k}]_{k=1}^{Nneg}$. The contrastive learning loss can be formulated as:
$\mathcal{L}_{NCE}=-log(\frac{e^{s_{p}}}{e^{s_{p}}+\sum_{k=1}^{Nneg}e^{s_{n}^{k}}})$ (4)4( 4 )
@@ IV-A Experimental Settings
To demonstrate the effectiveness of our method, we use two RGB-E datasets: VisEvent [8] and COESOT [4]. Details of the datasets can be found in their papers. We compare with some RGB-E trackers, including ViPT [10], CEUTrack [4], FENet [6], AFNet [7] and many other RGB-based trackers with two-modal input, e.g., SiamRPN++ [35], ATOM [36], STARK [39], MixFormer [20], etc. We adopt three metrics to evaluate trackers’ performance, including precision rate (PR), success rate (SR), and normalized precision rate (NPR).

### v3

@@ Abstract
The unique complementarity of frame-based and event cameras for high frame rate object tracking has recently inspired some research attempts to develop multi-modal fusion approaches. However, these methods directly fuse both modalities and thus ignore the environmental attributes, e.g., motion blur, illumination variance, occlusion, scale variation, etc. Meanwhile, insufficient interaction between search and template features makes distinguishing target objects and backgrounds difficult. As a result, performance degradation is induced especially in challenging conditions. This paper proposes a novel and effective Transformer-based event-guided tracking framework, called eMoE-Tracker, which achieves new SOTA performance under various conditions. Our key idea is to disentangle the environment into several learnable attributes to dynamically learn the attribute-specific features and strengthen the target information by improving the interaction between the target template and search regions. To achieve the goal, we first propose an environmental Mix-of-Experts (eMoE) module that is built upon the environmental Attributes Disentanglement to learn attribute-specific features and environmental Attributes Assembling to assemble the attribute-specific features by the learnable attribute scores dynamically. The eMoE module is a subtle router that prompt-tunes the transformer backbone more efficiently. We then introduce a contrastive relation modeling (CRM) module to emphasize target information by leveraging a contrastive learning strategy between the target template and search regions. Extensive experiments on diverse event-based benchmark datasets showcase the superior performance of our eMoE-Tracker compared to the prior arts. Project page: https://vlislab22.github.io/eMoE-Tracker/
@@
Due to the limited space in the main paper, we provide additional material for the proposed method and experimental results. Sec. VI introduce the datasets. Then, Sec. VII illustrates more details about the implementation and experiments. Afterward, we report more visual results and performance evaluation under different attributes in Sec. LABEL:sec:additional_results. In the end, Sec. IX makes summary for the complementary material.
@@ I INTRODUCTION
Figure 1: An illustration of the core idea of the environmental MoE (eMoE) module. This module acts as a subtle router to prompt-tune the frozen backbone encoder. The number of experts is determined by the attributes we decouple for the environmental conditions, and each expert is responsible for learning the attribute-specific features. All the learned features are assembled and added with the outputs from the backbone encoder at the corresponding layer for robust tracking representation.
@@
This has inspired research endeavors in developing event-guided, i.e., RGB-event (RGB-E) multi-modal tracking approaches [4, 5, 6, 7, 8, 9, 10]. These works can be divided into two categories based on the network structure: two-stream, i.e., siamese network [5, 6, 7] and one-stream trackers [4, 9, 10]. The former takes two identical branches to process RGB and event modalities separately. To better leverage the complementarity and increase the interaction between them, complex fusion modules are designed, thus leading to model complexity. The latter is usually based on the vision transformer (ViT) structure [11], where RGB and event tokens are concatenated and fed into the ViT backbone for feature encoding. Although they are free from the complex network structure, they fail to consider the impact of environmental attributes on tracking performance in challenging conditions. Meanwhile, inadequate interaction between search and template features makes distinguishing target objects and backgrounds difficult. Consequently, performance degradation is induced especially in challenging conditions. Intuitively, we raise a novel research question: how to design a one-stream framework that can distinguish the environmental attributes while enabling feature interaction for robust tracking under diverse visual conditions?
@@
In this paper, we propose a novel one-stream framework with an environmental Mixture-of-Experts structure (eMoE) along with a contrastive relation modeling (CRM) module to achieve robust tracking in challenging conditions, as shown in Fig. 1. The key insight is to disentangle the environmental attributes through learnable layers to dynamically learn the attribute-specific features for better tracking representation learning under challenging conditions. Specifically, to disentangle the environmental attributes, the eMoE module ((Sec. III-B)) is proposed to achieve two goals: (i) the environmental attributes disentanglement and (ii) the environmental attributes assembling. For the former, the eMoE module disentangles four attributes –illumination variance, motion blur, scale variance, and occlusion – to learn the attribute-specific features. This has been experimentally shown sufficient to reflect environmental effects on tracking given the advantages of event cameras (See Tab. IV). For the latter, each attribute-specific feature is assembled to build a more discriminative representation for tracking w.r.t.the attribute scores under the corresponding visual conditions, e.g., motion blur. For more efficient training, our eMoE module can be inserted into the arbitrary layers to prompt-tune the ViT backbone encoder. For the proposed CRM (Sec. III-B3) module, it aims to better distinguish the target features and background ones by introducing contrastive learning strategy. This subtly improves the interaction between the target template and search region and enhances the target objects. By integrating the eMoE and CRM modules, the output features can be more discriminative and less noisy for more robust tracking performance under diverse visual conditions.
@@ III Method
Figure 2: Overview of our proposed framework. The input of the whole network is the patch embeddings of RGB frames and stacked event frames. The concatenated two modal patches are fed into the backbone model and eMoE, and eMoE is inserted into the $l$-th layer of ViT to generate feature tokens which are combined with the tokens from the ViT encoder at the corresponding layer. $E^{l}$ is the ViT encoder at layer $l$. The CRM module gets the enhanced tokens to further improve the discriminability of target object.
@@ III-A1 Problem Setting
Given an initial target bounding box $B_{0}$ in a video, the goal of an RGB-based tracker is to learn a tracking model $T_{RGB}:\{I_{RGB},B_{0}\}\rightarrow B$ to estimate the bounding box in all subsequent frames $I_{RGB}$. In RGB-E tracking, event streams are introduced and stacked as event frames, extending the input to $(I_{RGB},I_{E})$, where the subscript $E$ indicates events. For details of representations for event data, we refer readers to [3, 29]. Therefore, the RGB-E tracking model can be represented as $T_{RGB-E}:\{I_{RGB},I_{E},B_{0}\}\rightarrow B$. We choose the transformer encoder and decoder of  [4] as our backbone encoder and decoder for better efficiency. The archetypical structure of the backbone encoder and decoder can be represented as $\mathcal{F}_{E}\circ\mathcal{F}_{D}$, where $\mathcal{F}_{E}:\{I_{RGB},I_{E},B_{0}\}\rightarrow\mathcal{T}_{RGB-E}$ denotes the backbone encoder and $\mathcal{F}_{D}:\mathcal{T}_{RGB-E}\rightarrow B$ represents the decoder which produces the estimated bounding box results $B$. The main body of $\mathcal{F}_{D}$ here is a vanilla vision transformer [11] containing 12 encoder layers. Each layer contains Multi-head Self-Attention (MSA), LayerNorm (LN), Feed-Forward Network (FFN) and residual connections. Before feeding inputs into the backbone network, RGB and event patches are projected into feature tokens adding with positional embeddings and then concatenated to RGB and event feature tokens $\mathcal{T}_{RGB-E}^{0}=[\mathcal{T}_{RGB}^{z},\mathcal{T}_{RGB}^{x},\mathcal{T}_{E}^{z},\mathcal{T}_{E}^{x}]$ as the inputs of the transformer encoder. Tokens through the $l$-th encoder layer $E^{l}$ can be represented as $\mathcal{T}_{RGB-E}^{l-1}$. The final layer encoder output is denoted as $\mathcal{T}_{RGB-E}^{L}$.
@@ III-B1 Overview
An overview of our eMoE-Tracker is shown in Fig. 2. The RGB and event inputs are first projected into a sequence of tokens and fed into backbone encoders and eMoE. The eMoE module aims to achieve environmental attributes disentanglement and environmental attributes assembling. The backbone encoder layers are frozen and the parameters are not updated. The eMoE module disentangles the environmental attributes to learn the attribute-specific and assembled features. Outputs from the eMoE module can be dynamically added to the tokens from the corresponding layer of the ViT backbone. Overall, the process can be formulated as follows:
@@
The eMoE module aims to achieve: i) environmental attributes disentanglement and ii) environmental attributes assembling. We now describe them.
i. Environmental Attributes Disentanglement. As shown in Fig. 3, to enable better learning of the attribute-specific features under various conditions, we manually annotate the visible-event datasets with four attribute labels including motion blur, illumination variance, scale variance, and occlusion at video level. Then, a mixture-of-experts network with four identical branches is designed to learn the attribute-specific features for each challenging condition. This allows us to capture more discriminative features and suppress the noises brought by other environmental attributes. All expert networks employ the CONV-MLP-CONV structure but with different parameters. Considering the $l$-th ViT layer, we assume that there are $K$ experts $\{f_{expert}^{l,i}(\mathcal{T}^{l}):\mathbb{R}^{N\times D}\rightarrow\mathbb{R}^{N\times D},i\in[1,K]\}$ to learn the attribute-specific features under the corresponding environmental condition, where $l$ denotes the layer of ViT backbone and $i$ represents the index of experts. Through each expert, we generate a series of attribute-specific features $\{\mathcal{H}_{i}^{l}\in\mathbb{R}^{N\times D},i\in[1,K]\}$ by expert $f_{expert}^{l,i}(\mathcal{T}^{l})$.
ii. Environmental Attribute Assembling. After obtaining the attribute-specific features under different decoupled environmental conditions, we should consider the different contributions of these features with the supervision of the ground truth attribute labels $G=[G_{1},G_{2},...,G_{K}]$. To achieve the goal, the assembling network employs the CONV-BatchNorm-ReLU-CONV-Sigmoid structure with two loops and is fed by all the RGB and event patch tokens to generate $K$ attribute scores for attribute-specific features, where $K$ denotes the number of experts. The learnable score indicates the ratio of different challenging types in the corresponding scenario, therefore the attribute-specific feature with a larger score should have a higher contribution to the assembling features to achieve the robust representation under various challenging conditions. Moreover, it can suppress the noise from other environmental attributes. Specifically, at $l$-th layer of the backbone, the attribute scores $W^{l,t}$ are generated from the assembling network $\{f_{a}^{l,t}(\mathcal{T}^{l}):\mathbb{R}^{N\times D}\rightarrow\mathbb{R}^{K}\}$, where $t$ represent the index of experts. The assembling feature $\mathcal{F}_{assemble}^{l}$ at layer $l$ can be formally calculated by $\sum_{t=1}^{K}W^{l,t}\mathcal{H}_{t}^{l}.$
@@ III-B3 Contrastive Relation Modeling
Apart from increasing the distinguishable ability of representations in different environmental conditions, we want to enhance the target information by introducing more interaction between target template and search region. Based on the fact that the features from the target template primarily contains target information, while the search regions include both the target and background information, we propose a CRM module by leveraging the contrastive learning strategy. As shown in Fig. 4, we first fuse two-modal patch tokens into fused target template tokens and search region tokens to better build the relationship. After fusion, we create positive pairs between features from target template and target object contents of search regions, and negative pairs between features from target template and background contents of search regions. To better strengthen the target object information, contrastive learning mechanism is exploited to pull the positive pairs closer and push negative pairs away thereby allowing the target object and background more distinguishable. The proposed CRM module effectively helps generate more unambiguous representation and achieve high performance in various challenging conditions.
@@ III-C Optimization
The main body of our RGB-E tracker $F_{RGB-E}$ is initialized by the transformer-based tracking backbone [22]. All the parameters $\theta$ we should update is only existing in eMoE and CRM. The optimization process can be formulated as
@@
where $L_{cls}$ is the focal loss [30] for classification, IoU loss [31] $L_{iou}$ and $L_{1}$ are exploited for bounding box regression, $\lambda_{iou}$ and $\lambda_{L_{1}}$ are regularization parameters. For more details, please refer to [22]. Additionally, we take the InfoNCE loss [32] as a contrastive learning loss for the CRM module. Given the fused target template tokens $\mathcal{T}_{fused}^{z}=[t_{fused}^{z,1},t_{fused}^{z,2},...,t_{fused}^{z,N_{z}}]$ from the ViT encoder, we compute the similarity $S=[s^{1},s^{2},...,s^{N_{x}}]$ between $\mathcal{T}_{fused}^{z}$ and the fused search region tokens $\mathcal{T}_{fused}^{x}=[t_{fused}^{x,1},t_{fused}^{x,2},...,t_{fused}^{x,N_{x}}]$. where $\tau$ is the temperature parameter and $s^{i}=sim(\mathcal{T}_{fused}^{z},t_{fused}^{x,i})/\tau$. Based on that, the search region tokens which contain the information inside the ground-truth bounding box can be selected as positive pairs including $N_{pos}$, where the similarity score is defined as $[s_{p}^{k}]_{k=1}^{N_{pos}}$. The left are negative pairs including $N_{neg}$, where the similarity score is set as $[s_{n}^{k}]_{k=1}^{N_{neg}}$. The contrastive learning loss can be formulated as:
$\mathcal{L}_{NCE}=-log(\frac{\sum_{k=1}^{N_{pos}}e^{s_{p}^{k}}}{\sum_{k=1}^{N_{pos}}e^{s_{p}^{k}}+\sum_{k=1}^{N_{neg}}e^{s_{n}^{k}}})$ (4)4( 4 )
@@ IV-A Experimental Settings
We use two datasets to demonstrate the effectiveness of our method: VisEvent [8] and COESOT [4]. We compare with some RGB-E trackers, including ViPT [10], CEUTrack [4], FENet [6], AFNet [7] and many other RGB-based trackers with two-modal input, e.g., SiamRPN++ [35], ATOM [36], STARK [39], MixFormer [20], etc. We adopt three metrics to evaluate trackers’ performance, including precision rate (PR), success rate (SR), and normalized precision rate (NPR).